OpenStudy (anonymous):

Emily brought some money on a trip to the aquarium. She used half of the money to buy an admission ticket. Then, she used eight dollars to buy lunch at the food court. She used three-fourths of what was left to buy a souvenir, leaving her with four dollars. How much money did Emily take to the aquarium?

OpenStudy (radar):

Don't you just love those worded problems. They check your ability to translate mathematical sentences to algebraic structures.

OpenStudy (radar):

Let x = the amount of money emily brought to the aquarium. We now sum up emily expenditures add the $4.00 she has left and we will have x or what loot she brought with her.

OpenStudy (radar):

Do you think that plan of attack will work?

OpenStudy (radar):

Answer please so I know you are there and awake.

OpenStudy (radar):

Then pay attention here, did you read the plan?

OpenStudy (anonymous):

yes i read it

OpenStudy (radar):

Good, we will now itemize Emily's expenses.....remember she brought x dollars. 1. Ticket cost 1/2 of what she brought with her. $ x/2 2. Lunch $8.00 we now total what she has left (x/2) - $8.00 The reason we did that is because spent 3/4 of what was left for a souvenir . 3. 3/4 (x/2 -$8.00)=(3/8)x-$6.00) that is the cost of the souvenir. 4 $4.00 is what she had left. Now if we sum up all of this it will equal what she brought.

OpenStudy (radar):

x=x/2 + $8.00 + (3/8) x - $6.00 + $4.00 x =(7/8)x + $6.00 (1/8)x=$6.00 x= $48.00 She brought $48.00 $48.00 $24.00 ticket $8.00 lunch $12.00 souvenir $4.00 left over

OpenStudy (radar):

Follow those steps and algebra you will get the same.